$80,000 is borrowed at 4% compounded monthly with regular monthly payments fully amortizing over 30 years. What is the monthly payment?
Answer #1
Monthly payment = Loan amount / Present value of annuity of 1
= $       80,000 / 209.4612
= $       381.93
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.00333)^-360)/0.00333 i = 4%/12 =            0.00333
=       209.4612 n = 30*12 = 360